# A Two-Loop Four-Point Form Factor at Function Level

### v2

@@ 1 Introduction
More recently, the four-point MHV form factor for the same ${\rm tr}\phi^{2}$ operator has been shown to exhibit antipodal self-duality on a four-dimensional (parity-preserving) slice of its five-dimensional phase space Dixon:2022xqh . This self-duality encompasses the previous duality, in the sense that the kinematic map in the four-point form factor relates two different limits, one which produces the three-point form factor, and the other produces the six-gluon scattering amplitude (because it is also a triple-collinear splitting amplitude). On the other hand, there is no fundamental understanding of why antipodal self-duality should exist, nor has it been verified yet beyond the symbol level.
@@
The three-point ${\rm tr}\phi^{2}$ form factor Dixon:2020bbt ; Dixon:2022rse has also been fixed at function level through eight loops. However, beyond one loop, the knowledge of the four-point MHV ${\rm tr}\phi^{2}$ form factor was, until very recently Guo:2024bsd , limited to symbol level Dixon:2022xqh and a special limit in which the operator has a light-like momentum Guo:2022qgv .111 The four-point ${\rm tr}\phi^{3}$ form factor, on the other hand, involves simpler Feynman integrals than ${\rm tr}\phi^{2}$; it has been computed at two loops using a bootstrap based on master integrals Guo:2021bym . (The three-point ${\rm tr}\phi^{3}$ form factor was very recently computed through six loops Basso:2024hlx ; Henn:2024pki .)
One difficulty in lifting the four-point MHV ${\rm tr}\phi^{2}$ form factor to a function is the large and intricate symbol alphabet: The alphabet for the three-point ${\rm tr}\phi^{2}$ form factor has only 6 letters, while that for the six-point and seven-point amplitude have 9 and 42 letters, respectively Goncharov:2010jf ; Golden:2013xva . The symbol alphabet for the four-point MHV ${\rm tr}\phi^{2}$ form factor at two (three) loops has 34 (88) letters Dixon:2022xqh . While seven-point scattering amplitudes have been successfully lifted from symbol-level to function level Dixon:2020cnr , the symbol alphabet in that case has only 42 letters, and is considerably easier to rationalize by a suitable choice of kinematical variables, compared with the 34 (88) letter alphabet for the four-point MHV ${\rm tr}\phi^{2}$ form factor.
Lifting the symbol of the four-point MHV ${\rm tr}\phi^{2}$ form factor to a full function would bring several new pieces of information. The numerical values, which are unavailable from only the symbol, could (if evaluated to high enough loop order) probe the radius of convergence of perturbation theory Dixon:2014voa ; Dixon:2015iva ; Caron-Huot:2019vjl . It might also be possible to interpolate or extrapolate to strong coupling where a minimal surface formulation is available Alday:2007hr ; Alday:2009dv ; Maldacena:2010kp . In the case of amplitudes, such interpolation is relatively simple at special kinematic points called origins Basso:2020xts ; Basso:2022ruw , but it is also possible in the OPE limit Basso:2013vsa ; Sever:2020jjx . Antipodal self-duality can be analyzed beyond symbol level. Various Minkowski factorization limits can also be studied, such as multi-Regge kinematics, the self-crossing (or pseudo-double-parton scattering) limit Georgiou:2009mp ; Dixon:2016epj ; Caron-Huot:2019vjl , and the light-like limit Guo:2022qgv .
@@
Very recently, the four-point MHV ${\rm tr}\phi^{2}$ form factor has also been computed at two loops Guo:2024bsd at function level, by using unitarity cut methods to obtain the loop integrands in $D$ spacetime dimensions. Integration-by-parts reduction was then used to write the result in terms of the basis of two-loop, non-planar five-point master integrals with one external mass provided in ref. Abreu:2023rco . These master integrals are provided in a $2\to 3$ scattering configuration. Ref. Guo:2024bsd also used AMFlow Liu:2022chg to compute a couple of numerical values in (pseudo-)Euclidean or $1\to 4$ decay kinematics. As we focus mainly on the Euclidean region in our paper, and provide more analytic results, our results are complementary to those of ref. Guo:2024bsd .
This paper is organized as follows. In Sec. 2 we introduce our notation and review basic properties of form factors and multiple polylogarithms. In Sec. 3 we define the space of functions used to describe the four-point MHV ${\rm tr}\phi^{2}$ form factor. We also describe the three-parameter rational surface kinematics. We then fix all the beyond-the-symbol constants by utilizing the available information. We provide the result for the remainder function in the bulk in Sec. 4. We conclude in Sec. 5 with a discussion of future research directions enabled by this work.
Many of the explicit results of this paper are rather lengthy, so they are included as computer-readable ancillary files: AntipodalAlphabet.m specifies the symbol alphabet we work with. R42funcCoTable.m gives the iterated coproducts of the function in terms of a set of independent lower-weight functions we call $P$ functions. R42_rational.m is the representation of the remainder function on the rational surface in terms of multiple polylogarithms. Prat.m provides the same representation for the $P$ functions. PTT2to0_xy.m provides the $P$ functions in the OPE limit (defined in Sec. 3). R42_OPE.txt provides the set of coefficient functions describing the remainder function in the OPE limit (see Appendix A).
@@ 2.1 BDS-like normalized form factors
In this paper, we study the MHV form factor for the chiral stress energy tensor in planar $\mathcal{N}=4$ SYM. One representative for this BPS-protected operator super-multiplet is ${\rm tr}\phi^{2}$, where $\phi^{2}$ is some traceless (non-Konishi) scalar bilinear, and the trace “${\rm tr}$” is over the large-$N_{c}$ $SU(N_{c})$ gauge group. One component of the super four-point form factor is the matrix element of ${\rm tr}\phi^{2}$ with two massless scalars and two same-helicity gluons:
@@
Here “$+$ cyclic” means to add the three images under the cyclic transformation ${\cal C}$ in eq. (7).
@@
$\hskip 0.0pt\hbox{}-\ln(v_{1}v_{4}-u_{1})[2\ln v_{4}-2\ln(1-v_{4})+\ln u_{1}+\ln u_{2}+\ln u_{3}-\ln u_{4}]$
$\hskip 0.0pt\hbox{}-2\ln(1-v_{4})[2\ln v_{4}+\ln u_{3}]+2\ln v_{4}[\ln u_{2}+\ln u_{3}-\ln u_{4}]$
$\hskip 0.0pt\hbox{}+\ln u_{1}\ln u_{2}-\ln u_{2}\ln u_{4}+\ln u_{4}\ln u_{1}+2\zeta_{2}\,.$
@@ 2.2 Polylogarithms and (antipodable-)symbol alphabet
where the sum is over letters $\phi$ which belong to the symbol alphabet $\Phi$, and the $\phi$-coproducts $F^{\phi}$ appearing in eq. (18) have weight $n-1$. In integral form, MPLs are commonly expressed as $G$ functions,
@@ 3.1 Function space
The framework for integrating up the remainder function from lower-weight functions is the coproduct formalism Dixon:2013eka ; Caron-Huot:2019bsq . We construct a set of basis functions, $\{F^{(w)}_{i}\}$, for each weight $w$ up to $2L$. The sets have dimensions $|F^{(w)}|$ and are big enough to contain all the (multiple) coproducts of the form factor or amplitudes. The different bases are linked to each other by the coproducts $\Delta_{w-1,1}$, which connect two consecutive bases, for weights $w-1$ and $w$, via a three-index tensor $T^{(w)}_{ij\phi}$:
@@
Here $i,j$ are indices labelling the basis functions at weight $w$ and $w-1$, respectively; $\phi$ are letters in the symbol alphabet $\Phi$; $T^{(w)}_{ij\phi}$ are rational numbers filling out a three-index tensor with dimension $|F^{(w)}|\times|F^{(w-1)}|\times|\Phi|$. The symbol-level information corresponds to the coefficients of functions with nonvanishing symbols, whereas the coefficients of MZVs within the $\{F^{(w)}_{i}\}$ are yet to be fixed. Generally the dimensions $|F^{(w)}|$ have to increase from the symbol-level version, in order to accommodate the MZVs. In our two-loop case, the increase will be very modest, just an increase of one function at weight 2, to account for $\zeta_{2}$.
@@ 3.2.1 Rational surface and soft/collinear factorization
is mapped to itself through the reflection (31) which exchanges $v_{1}\leftrightarrow v_{2}$. In fact, the symbol of $\mathcal{R}^{(2)}_{4}(u_{3},v_{1},v_{2})$ vanishes at the rational surface boundary $u_{3}\to 0$. We find, after implementing other constraints, that $\mathcal{R}^{(2)}_{4}(u_{3},v_{1},v_{2})$ also vanishes at function level for $u_{3}\to 0$. So the way the dihedral symmetry of $\mathcal{R}^{(2)}_{4}$ is implemented on this boundary is rather trivial, $0\leftrightarrow 0$.
@@
$\lim_{\hat{v}\to 1,\hat{w}\to 0}R^{(2)}_{6}(\hat{u},\hat{v},\hat{w})$ $=$ $\ln\hat{w}\bigl{[}2{\rm Li}_{3}(\hat{u})-\ln\hat{u}{\rm Li}_{2}(\hat{u})\bigr{]}-4{\rm Li}_{4}(\hat{u})+2{\rm Li}_{4}\left(\frac{-\hat{u}}{1-\hat{u}}\right)-\frac{1}{2}\bigl{[}{\rm Li}_{2}(\hat{u})\bigr{]}^{2}$ (37)
$\hskip 0.0pt\hbox{}+\ln\hat{u}\biggl{[}2{\rm Li}_{3}(\hat{u})-\ln(1-\hat{u}){\rm Li}_{2}(\hat{u})-\frac{1}{3}\ln^{3}(1-\hat{u})\biggr{]}+\frac{1}{12}\ln^{4}(1-\hat{u})$
$\hskip 0.0pt\hbox{}+\zeta_{2}\bigl{[}2{\rm Li}_{2}(\hat{u})+\ln^{2}(1-\hat{u})\bigr{]}\,,$
@@
In summary, the triple-collinear boundary condition is $\mathcal{R}_{4}^{(2)}(u_{3},v_{1},v_{2})|_{v_{1}\to 0}=R_{6}^{(2)}(\hat{u},1,\hat{w})$, where $\hat{u},\hat{w}$ are given in eq. (38).
@@ 3.2.2 FFOPE limit: Near multi-collinear factorization
and this is linear in $X$. For instance, take the two letters $X+Y-XY$ and $X^{2}+(1-X^{2})Y$. They contain $X^{2}$ but only contribute to $\Phi^{(Y)}_{XY}$ as $(1-X)X^{2}-X(1-X^{2})=-X(1-X)$.
@@ 4.2 Antipodal duality beyond the symbol
${\rm disc}_{u_{2}}\Bigl{[}\mathcal{R}^{(2)}_{4}(\eta)|_{\rm finite\,part}\Bigr{]}$ =±$=\pm i\pi\Bigl{[}4H_{-1,-1,0}(\eta)-12H_{-1,0,0}(\eta)-4H_{-1,1,0}(\eta)-8H_{0,-1,0}(\eta)$ (56)
$+24H_{0,0,0}(\eta)+8H_{0,1,0}(\eta)-4H_{1,-1,0}(\eta)+12H_{1,0,0}(\eta)+4H_{1,1,0}(\eta)\Bigr{]}\,.$
@@
The symbol alphabet on the $u_{1},u_{2},u_{3}\to 0$, $10<v_{1},v_{2}<$ 1 surface, which contains the green and red segments, simplifies to $\{v_{1},v_{2},1-v_{1},1-v_{2},1-v_{1}-v_{2},v_{1}+v_{2}\}$. The finite part of the remainder function can be represented on this surface in terms of $G$-functions, as shown in Appendix C. Near $v_{2}\to 1$, the expression becomes simple logarithms,
@@ 5 Conclusion
The Wilson loops dual to scattering amplitudes and form factors develop singularities when the Wilson lines intersect. The scattering kinematics are similar to multi-parton scattering in hadronic collisions, but here there are only two incoming gluons, each of which splits into two almost on-shell collinear virtual partons that then scatter off each other Dixon:2016epj . (In the case of the four-point form factor, one pair annihilates into an operator.) For the case of six-gluon scattering, the self-crossing limit is characterized by letting one cross ratio $\hat{u}\to 1$, while the other two cross ratios are set equal to each other, $\hat{w}=\hat{v}$. The four-point form factor self-crossing kinematics constrain three of the five kinematic variables. They are a bit simpler to describe in the OPE variables [^2] We thank Benjamin Basso for a discussion on this point.: $F_{2}=1$, and $T^{2}=-S^{2}(1+T_{2}^{2}+S_{2}T_{2})/(S_{2}T_{2})$. The logarithmic divergences as one approaches the self-crossing limit can be characterized by a renormalization group equation Korchemsky:1993hr ; Korchemskaya:1994qp , $\mu\frac{\partial}{\partial\mu}\mathcal{W}_{i}=-\Gamma^{ij}(\gamma,g)\mathcal{W}_{j}$, where $\mathcal{W}_{1}$, $\mathcal{W}_{2}$ are the framed Wilson loops with crossing and disconnected topology Dixon:2016epj , which mix with each other under renormalization. The renormalization scale $\log\mu$ can be chosen as $\log(1-\hat{u})$ in the six-gluon case. The cross anomalous dimension matrix $\Gamma^{ij}$ has kinematic dependence only on the local crossing angle $\gamma$, and can be determined by comparing with the six-gluon remainder function Dixon:2016epj , or using an all-orders formula based on integrability Caron-Huot:2019vjl . For the four-gluon form factor, we also expect a similar evolution of self-crossing singularities, but again the details remain to be investigated.
@@ Acknowledgments
We thank Benjamin Basso, Ömer Gürdoğan, Zhengjie Li, Yu-Ting Liu, Andrew Mcleod, Matthias Wilhelm, and Gang Yang for useful discussions. We are grateful to the Simons Center for Geometry and Physics for hospitality during the program “Solving ${\cal N}=4$ super-Yang-Mills theory via Scattering Amplitudes”. LD also thanks Humboldt University Berlin for hospitality. This research was supported by the US Department of Energy under contract DE–AC02–76SF00515.
@@ Appendix B Remainder function in 2D kinematics
$+4\Bigl{[}G_{1}(v_{1})\Bigl{(}G_{0}(v_{1})-G_{0}(v_{2})\Bigr{)}-2G_{0,1}(1-v_{1})\Bigr{]}\biggl{[}G_{0,1}\Bigl{(}\frac{-v_{2}}{v_{1}-v_{2}}\Bigr{)}-G_{0,1}\Bigl{(}\frac{-v_{2}}{1-v_{1}-v_{2}}\Bigr{)}\biggr{]}$
@@
$+24G_{1}(v_{2})\Bigl{(}G_{0,0,1}(1-v_{1})+G_{0,0,1}(v_{1})\Bigr{)}+12\Bigl{(}G_{0}(v_{1})+G_{1}(v_{1})+G_{0}(v_{2})\Bigr{)}G_{0,0,1}(1-v_{2})$
$+4G_{1}(v_{2})G_{0,0,1}(1-v_{2})+12\Bigl{(}G_{0}(v_{1})+G_{1}(v_{1})+G_{1}(v_{2})\Bigr{)}G_{0,0,1}(v_{2})+4G_{0}(v_{2})G_{0,0,1}(v_{2})$

### v3

@@ 1 Introduction
More recently, the four-point form factor for the same ${\rm tr}\phi^{2}$ operator has been shown to exhibit antipodal self-duality on a four-dimensional (parity-preserving) slice of its five-dimensional phase space Dixon:2022xqh . This self-duality encompasses the previous duality, in the sense that the kinematic map in the four-point form factor relates two different limits, one which produces the three-point form factor, and the other produces the six-gluon scattering amplitude (because it is also a triple-collinear splitting amplitude). On the other hand, there is no fundamental understanding of why antipodal self-duality should exist, nor has it been verified yet beyond the symbol level.
@@
The three-point ${\rm tr}\phi^{2}$ form factor Dixon:2020bbt ; Dixon:2022rse has also been fixed at function level through eight loops. However, beyond one loop, the knowledge of the four-point ${\rm tr}\phi^{2}$ form factor is presently limited to symbol level Dixon:2022xqh and a special limit in which the operator has a light-like momentum Guo:2022qgv . The four-point ${\rm tr}\phi^{3}$ form factor, on the other hand, involves simpler Feynman integrals than ${\rm tr}\phi^{2}$; it has been computed at two loops using a bootstrap based on master integrals Guo:2021bym . (The three-point ${\rm tr}\phi^{3}$ form factor was very recently computed through six loops Basso:2024hlx ; Henn:2024pki .)
One difficulty in lifting the four-point ${\rm tr}\phi^{2}$ form factor to a function is the large and intricate symbol alphabet: The alphabet for the three-point ${\rm tr}\phi^{2}$ form factor has only 6 letters, while that for the six-point and seven-point amplitude have 9 and 42 letters, respectively Goncharov:2010jf ; Golden:2013xva . The symbol alphabet for the ${\rm tr}\phi^{2}$ form factor at two (three) loops has 34 (88) letters Dixon:2022xqh . The latter is considerably larger, and harder to rationalize, than the seven-point amplitude’s alphabet.
Lifting the symbol of the four-point form factor to a full function would bring several new pieces of information. The numerical values, which are unavailable from only the symbol, could (if evaluated to high enough loop order) probe the radius of convergence of perturbation theory Dixon:2014voa ; Dixon:2015iva ; Caron-Huot:2019vjl . It might also be possible to interpolate or extrapolate to strong coupling where a minimal surface formulation is available Alday:2007hr ; Alday:2009dv ; Maldacena:2010kp . In the case of amplitudes, such interpolation is relatively simple at special kinematic points called origins Basso:2020xts ; Basso:2022ruw , but it is also possible in the OPE limit Basso:2013vsa ; Sever:2020jjx . Antipodal self-duality can be analyzed beyond symbol level. Various Minkowski factorization limits can also be studied, such as multi-Regge kinematics, the self-crossing (or pseudo-double-parton scattering) limit Georgiou:2009mp ; Dixon:2016epj ; Caron-Huot:2019vjl , and the light-like limit Guo:2022qgv .
@@
Very recently, the four-point ${\rm tr}\phi^{2}$ form factor has also been computed at two loops Guo:2024bsd at function level, by using unitarity cut methods to obtain the loop integrands in $D$ spacetime dimensions. Integration-by-parts reduction was then used to write the result in terms of the basis of two-loop, non-planar five-point master integrals with one external mass provided in ref. Abreu:2023rco . These master integrals are provided in a $2\to 3$ scattering configuration. Ref. Guo:2024bsd also used AMFlow Liu:2022chg to compute a couple of numerical values in (pseudo-)Euclidean or $1\to 4$ decay kinematics. As we focus mainly on the Euclidean region in our paper, our results are complementary to those of ref. Guo:2024bsd .
This paper is organized as follows. In Sec. 2 we introduce our notation and review basic properties of form factors and multiple polylogarithms. In Sec. 3 we define the space of functions used to describe the form factor, and the three-parameter rational surface kinematics. We then fix all the beyond-the-symbol constants by utilizing the available information. We provide the result for the remainder function in the bulk in Sec. 4. We conclude in Sec. 5 with a discussion of future research directions enabled by this work.
Many of the explicit results of this paper are rather lengthy, so they are included as computer-readable ancillary files: AntipodalAlphabet.m specifies the symbol alphabet we work with. R42funcCoTable.m gives the iterated coproducts of the function in terms of a set of independent lower-weight functions we call $P$ functions. R42_rational.m is the representation of the remainder function on the rational surface in terms of multiple polylogarithms. Prat.m provides the same representation for the $P$ functions. PTT2to0_xy.m provides the $P$ functions in the OPE limit (defined in Sec. 3). R42_OPE.txt provides the set of coefficient functions describing the remainder function in the OPE limit (see Appendix A). R42_2d.m provides the remainder function on the 2D kinematics (see Appendix B).
@@ 2.1 BDS-like normalized form factors
In this paper, we study the MHV form factor for the chiral stress energy tensor in planar $\mathcal{N}=4$ SYM. One representative for this BPS-protected operator super-multiplet is ${\rm tr}\phi^{2}$, where $\phi^{2}$ is some traceless (non-Konishi) scalar bilinear, and the trace $\rm tr$}” is over the large-$N_{c}$ $SU(N_{c})$ gauge group. One component of the super four-point form factor is the matrix element of ${\rm tr}\phi^{2}$ with two massless scalars and two same-helicity gluons:
@@
Here “++ cyclic” means to add the three images under the cyclic transformation $\cal C$} in eq. (7).
@@
$\thinspace\hbox{}-\ln(v_{1}v_{4}-u_{1})[2\ln v_{4}-2\ln(1-v_{4})+\ln u_{1}+\ln u_{2}+\ln u_{3}-\ln u_{4}]$
$\thinspace\hbox{}-2\ln(1-v_{4})[2\ln v_{4}+\ln u_{3}]+2\ln v_{4}[\ln u_{2}+\ln u_{3}-\ln u_{4}]$
$\thinspace\hbox{}+\ln u_{1}\ln u_{2}-\ln u_{2}\ln u_{4}+\ln u_{4}\ln u_{1}+2\zeta_{2}\,.$
@@ 2.2 Polylogarithms and (antipodable-)symbol alphabet
where the sum is over letters $\phi$ which belong to the symbol alphabet $\Phi$, and the $\phi$-coproducts $F^{\phi}$ appearing in eq. (18) have weight n1n-1. In integral form, MPLs are commonly expressed as $G$ functions,
@@ 3.1 Function space
The framework for integrating up the remainder function from lower-weight functions is the coproduct formalism Dixon:2013eka ; Caron-Huot:2019bsq . We construct a set of basis functions, $\{F^{(w)}_{i}\}$, for each weight $w$ up to $2L$. The sets have dimensions $|F^{(w)}|$ and are big enough to contain all the (multiple) coproducts of the form factor or amplitudes. The different bases are linked to each other by the coproducts $\Delta_{w-1,1}$, which connect two consecutive bases, for weights w1w-1 and $w$, via a three-index tensor $T^{(w)}_{ij\phi}$:
@@
Here $i,j$ are indices labelling the basis functions at weight $w$ and w1w-1, respectively; $\phi$ are letters in the symbol alphabet $\Phi$; $T^{(w)}_{ij\phi}$ are rational numbers filling out a three-index tensor with dimension $|F^{(w)}|\times|F^{(w-1)}|\times|\Phi|$. The symbol-level information corresponds to the coefficients of functions with nonvanishing symbols, whereas the coefficients of MZVs within the $\{F^{(w)}_{i}\}$ are yet to be fixed. Generally the dimensions $|F^{(w)}|$ have to increase from the symbol-level version, in order to accommodate the MZVs. In our two-loop case, the increase will be very modest, just an increase of one function at weight 2, to account for $\zeta_{2}$.
@@ 3.2.1 Rational surface and soft/collinear factorization
is mapped to itself through the reflection (31) which exchanges $v_{1}\leftrightarrow v_{2}$. This symmetry requires $\mathcal{R}^{(2)}_{4}(u_{3},v_{1},v_{2})$ to vanish at the rational surface boundary $u_{3}\to 0$.
@@
[…]
@@
[…]
@@ 3.2.2 FFOPE limit: Near multi-collinear factorization
and this is linear in $X$. For instance, take the two letters X+YXYX+Y-XY and $X^{2}+(1-X^{2})Y$. They contain $X^{2}$ but only contribute to $\Phi^{(Y)}_{XY}$ as $(1-X)X^{2}-X(1-X^{2})=-X(1-X)$.
@@ 4.2 Antipodal duality beyond the symbol
[…]
@@
The symbol alphabet on the $u_{1},u_{2},u_{3}\to 0$, $0<v_{1},v_{2}<1$ surface, which contains the green and red segments, simplifies to $\{v_{1},v_{2},1-v_{1},1-v_{2},1-v_{1}-v_{2},v_{1}+v_{2}\}$. The finite part of the remainder function can be represented on this surface in terms of $G$-functions, as shown in Appendix C. Near $v_{2}\to 1$, the expression becomes simple logarithms,
@@ 5 Conclusion
[…]
@@ Acknowledgments
We thank Benjamin Basso, Ömer Gürdoğan, Zhengjie Li, Yu-Ting Liu, Andrew Mcleod, Matthias Wilhelm, and Gang Yang for useful discussions. We thank Zhenjie Li and Song He for pointing out the error in Eq. (B.1). We are grateful to the Simons Center for Geometry and Physics for hospitality during the program “Solving ${\cal N}=4$ super-Yang-Mills theory via Scattering Amplitudes”. LD also thanks Humboldt University Berlin for hospitality. This research was supported by the US Department of Energy under contract DE–AC02–76SF00515.
@@ Appendix B Remainder function in 2D kinematics
$+4\Bigl[G_{1}(v_{1})\Bigl(G_{0}(v_{1})-G_{0}(v_{2})\Bigr)-2G_{0,1}(1-v_{1})\Bigr]G_{0,1}\Bigl(\frac{-v_{2}}{v_{1}-v_{2}}\Bigr)$
$-4\Bigl[G_{0}(v_{1})\Bigl(G_{1}(v_{1})+G_{0}(v_{2})\Bigr)-2G_{0,1}(1-v_{1})\Bigr]G_{0,1}\Bigl(\frac{-v_{2}}{1-v_{1}-v_{2}}\Bigr)$
@@
[…]
@@
The function is also available in Mathematica format in the ancillary file R42_2d.m.